OpenStudy (tmagloire1):

What is the area of the largest rectangle with lower base on the x-axis and upper vertices on the curve y = 3 − x^2?

OpenStudy (anonymous):

Okay here is the answer

OpenStudy (anonymous):

Area = x * y We need to find our X, because the height is a function of X. You can divide this problem into 2 rectangles, each one to both sides of the y axis, and the "big" rectangle will be the sum of both. so, the rectangle will have area = x * (12 - x^2) you have to maximize it, so we differentiate it: d(area)/dx = 12-3*x^2 = 0. then x = 2 the area would be 2*(12-4) = 16 lets see if its maximum or minimum... if we choose x = 0, we would get a height o 12, but an area of 0, so its a maximum (without doing any derivative tests :D ) So the area of both rectangles would be 32 units squared

OpenStudy (tmagloire1):

._. that wasn't even my problem. where did you get those numbers from

OpenStudy (tmagloire1):

but ill try this way and see what i get

OpenStudy (amistre64):

they used a different curve, but the process seems good enough.

OpenStudy (tmagloire1):

i got 4! thanks for the process